OpenStudy (anonymous):

3 + x = 4 is the same as 4 = 3 + x. This is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

Associative Property of Addition

OpenStudy (anonymous):

An equation is shown below: (4x - 1) + (2x + 2) = 1 Which statement shows a correct next step in solving the equation? The equation can become 6x + 1 = 1 by combining like terms. The equation can become 3x + 4x = 1 by solving the terms in the brackets. The equation can become 3x + 4x = 1 by applying the associative property of multiplication. The equation can become 6x + 1 = 1 by applying the distributive property.

OpenStudy (anonymous):

The equation can become 3x + 4x = 1 by applying the associative property of multiplication.

OpenStudy (yanasidlinskiy):

i think it would be c

OpenStudy (anonymous):

Evaluate: -6 - (-1) -7 5 -5 6

OpenStudy (yanasidlinskiy):

-5

OpenStudy (anonymous):

thnks

OpenStudy (yanasidlinskiy):

np:)

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Which of the following is the most appropriate unit to describe the rate at which data is transferred using the internet? Megabits per second, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits. Seconds per megabit, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits. Seconds per megabit, because the independent quantity is amount of data in megabits and dependent quantity is time in seconds. Megabits per second, because the independent quantity is amount of data in megabits and dependent quantity is time in seconds.

OpenStudy (anonymous):

Megabits per second, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits.

OpenStudy (anonymous):

Jessica is baking a cake. The recipe says that she has to mix 96 grams of sugar to the flour. Jessica knows that 1 cup of this particular sugar has a mass of 128 grams. She added of a cup of sugar to the flour. Should Jessica add more sugar to make the exact recipe or did she go over and by what amount? She went over by of a cup She needs to add of a cup She went over by of a cup She needs to add of a cup

OpenStudy (anonymous):

yall a real help bro

OpenStudy (anonymous):

well if she only added 1/2 cups, that means she only added (128/2) grams of sugar, which is 64 grams. From here you can see that she still needs more sugar so 1 and 3 are already out of the question. Now to find out how much more she needs to add, it's 32 grams. so 32/128= 1/4 and that should give you your answer **96g-64g=32 grams**

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Edward wants to determine which skateboarder won a race. The actual time taken by four skateboarders to skate the same distance is shown below. Skateboarder A B C D Time (in seconds) 4.28 4.24 4.26 4.25 If the stopwatch can measure to the nearest tenth of a second, which skateboarder will show as having a different time from the others? Skateboarder A Skateboarder B Skateboarder C Skateboarder D

OpenStudy (anonymous):

umm.. the first question u asked wouldn't be Associative....it would be Commutative Property of Addition

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

how

OpenStudy (anonymous):

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115

OpenStudy (anonymous):

u are just changing the order in which the numbers are.... the Commutative Property of Addition is : a+b = b+a

OpenStudy (anonymous):

Well we know that the total for one day was $115, so that doesn't change. Also it says $40 per day, and the question states it was for one day so that stays the same.

OpenStudy (anonymous):

okay so its b

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

OpenStudy (anonymous):

so its 40x+15=115

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

no....40 + 15x = 115

OpenStudy (anonymous):

thnks @ crabby_patty

OpenStudy (anonymous):

The length of the base of a triangle is 8.5 cm and its height is 5.4 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. (Round your answer to the appropriate significant digits)

OpenStudy (anonymous):

cause if u were to solve it, x =5 which is the cost per extra mile...............and ur welcome

OpenStudy (anonymous):

sry hate those kind of problems.....can't help

OpenStudy (anonymous):

A car repair shop charged $80 per hour for labor and $210 for the parts of a car. If the total cost of repair for the car was $490, how many hours did it take to repair the car? 0.59 2.71 3.50 8.75 -------------------------------------------------------------------------------- Question 16 (Multiple Choice Worth 1 points) (01.02) Joseph borrowed a book from a library. The library charged a fixed rental for the book and a late fee for every day the book was overdue. The expression below shows the charges Joseph paid for the book when he returned it x days after the due date. 2 + 0.25x What does the coefficient of the expression represent? The fixed rental for the book The total late fee for the book The late fee per day for the book The number of days the book was overdue -------------------------------------------------------------------------------- Question 17 (Multiple Choice Worth 1 points) (01.02) A pet shop had 8 identical fish tanks in the shape of rectangular prisms. The width and height of each tank was (x - 7) inches and the length was 8 inches more than its width. The expression below shows the maximum amount of water which all the tanks together can hold. 8(x + 1)(x - 7)2 What does the factor 8(x + 1) represent? The total height of all the tanks The total length of all the tanks The total volume of all the tanks The total base area of all the tanks -------------------------------------------------------------------------------- Question 18 (Multiple Choice Worth 1 points) (01.02) x • x = ? 2x x2 x(x + 2) x + 2 -------------------------------------------------------------------------------- Question 19 (Multiple Choice Worth 1 points) (01.02) Simplify the following expression. 2x - 8y + 3x2 + 7y - 12x. 3x2 - 14x - y 3x2 - 10x - y 3x2 - 10x - 15y 3x2 + 10x - y -------------------------------------------------------------------------------- Question 20 (Multiple Choice Worth 1 points) (01.01) Evaluate: [(21 + 6) - 32] ÷ 9 • 2 4 4.6 1 52 -------------------------------------------------------------------------------- Question 21 (Multiple Choice Worth 1 points) (01.01) Evaluate: 32 + (8 - 2) • 4 - 30 58 28 31

OpenStudy (anonymous):

u there